OpenStudy (anonymous):

Find the last three digits of the Number \(2003^{2002^{2001}}\).

OpenStudy (kinggeorge):

This is assuming \(2003^{(2002^{2001})}\) To find the last three digits, we want to take the number mod 1000. Thus, the question is equivalent to asking what is\[3^{2002^{2001}} \mod 1000\]Also, we know that\(\phi(1000)=400\) So \[3^{400} \equiv 1 \mod 1000\]. Now, we need to find \(2002^{2001} \mod 400\) Using the same method, we have that it's equivalent to \(2^{2001} \mod 400\), and \(\phi(400)=160\). Thus, \(2001 = 12*160 + 81\), so we need to calculate \(2^{81} \mod 400\). Using successive squaring this is relatively easy to do by hand, and you find that \(2^81 \equiv 352 \mod 400\) Finally, we need to calculate \(3^{352} \mod 1000\). Once again, we can use successive squaring, and we find that \(3^{352} \equiv 241 \mod 1000\).

OpenStudy (kinggeorge):

If we're assuming \((2003^{2002})^{2001}\), the answer is different.

OpenStudy (anonymous):

i am going to make a guess that it is the second one

OpenStudy (anonymous):

\[2003^{400}\equiv1 (\text{mod} 1000)\] i think

OpenStudy (kinggeorge):

If it's the second one, it's almost trivial. \(2002*2001 = 4006002 = 400k+2\) for some \(k\), so it's \(3^2 \equiv 9 \mod 1000\).

OpenStudy (anonymous):

no that is wrong what i wrote

OpenStudy (anonymous):

it is 9

OpenStudy (kinggeorge):

Your first statement was correct, since the gcd of 2003 and 1000 is 1, \(2003^{400} \equiv 1 \mod 1000\).

OpenStudy (anonymous):

and 9 only leaves remainders of 1 and 9, so answer is 9 if i am thinking clearly

OpenStudy (anonymous):

@KingGeorge correct me if i am wrong, but we can use \[a^{\phi(1000)}\equiv 1 (\text{mod }1000)\] yes?

OpenStudy (kinggeorge):

Correct.

OpenStudy (kinggeorge):

If \(gcd(a, 1000) =1 \)

OpenStudy (anonymous):

ah right. guess we need that too, i forgot. but it is true in any case

OpenStudy (anonymous):

i mean in this case

OpenStudy (kinggeorge):

It is true in this case. More formally,\[(2003^{2002})^{2001} \equiv 2003^{4006002}\equiv 3^{10015*400}\cdot 3^2 \mod 1000\]So \[3^{10015*400}\cdot 3^2 \equiv 1*3^2 \equiv 9 \mod 1000\]

OpenStudy (anonymous):

short and sweet